Rolf Steps is the production manager for a local manufacturing firm. This company produces staplers and other items. The holding cost is $2 per unit per year. The cost of setting up the production line for this is $25. There are 200 working days per year. The production rate for this product is 80 per day. If the production order quantity is 200 units, what was the daily demand (rounded to the nearest whole unit)?

6 units

7 units

8 units

9 units

None of these
Answer #1

8

EOQ(Economic Order Quantity) = sqroot{(2*Annual Demand*setup cost)/(Holding cost)}

200 = sqroot{(2*Annual Demand*25)/(2)}

Squaring Both sides.

200*200 = (2*Annual Demand*25)/(2)

Annual Demand = (200*200*2)/(50)

Annual Demand = 1600

Daily Demand = Annual Demand/Number of working days

= (1600/200)

= 8